OpenStudy (anonymous):

If I were to simply (3x^3)^2 would the correct answer be 9x^5. I can't remember if I use addition on the exponents on the top.

OpenStudy (anteater):

You would use addition if it were x^2 times x^3. For (x^2)^3 you would multiply the exponents.

OpenStudy (anteater):

To prove this to yourself: (x^2)^3 = (x^2)(x^2)(x^2) = x^(2+2+2)

OpenStudy (anteater):

So, 9x^6 Hope that helps with future problems. :)

OpenStudy (anonymous):

Thanks I just wasn't sure. but that helps a lot thanks.

OpenStudy (anteater):

Welcome :)